# StateFlow: Enhancing LLM Task-Solving through State-Driven Workflows

### v2

@@ 1 Introduction
LLMs have increasingly been employed to solve complex, multi-step tasks. Specifically, they have been applied to tasks that require interactions with environments (Yang et al., 2023a; Yao et al., 2022a; Shridhar et al., 2020) and those tasks that can benefit from utilizing tools such as web search and code execution (Mialon et al., 2023; Wu et al., 2023b; Wang et al., 2023c). In approaching these tasks, there is typically a desired workflow, or plan of actions based on heuristics that could improve the efficiency of task solving (Kim et al., 2023; Wu et al., 2023a). A common practice in the context of LLMs, such as ReAct (Yao et al., 2022b) and the vast customization of GPTs, is to write a single prompt that instructs the models to follow a desired procedure to solve the task. The LLM is called iteratively with the same instruction, along with previous actions and feedback from tools/environments. This relies on LLMs’ innate capability to determine the current task-solving status and perform subsequent actions autonomously. Despite the impressive abilities of LLMs, it is still unrealistic to expect LLMs to always make the correct judgment of the status of current progress. It is also almost impossible to reliably track these judgments and their decisions of subsequent action trajectory. Given these considerations, we pose the research question: How can we exert more precise control and guidance over LLMs?
@@ 4.1 InterCode Benchmark
[…]
@@
[…]
@@
SR Turns Error Cost
% $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$
StateFlow 63.73 5.67 6.82 3.82
No_Verify 62.28 5.18 5.96 3.68
No_Error 58.80 5.72 11.6 4.05
No_Obsrve 57.83 6.00 17.0 4.64
@@ 4.2 ALFWorld
[…]
@@ 5 Related work
Different prompting frameworks have been proposed to enhance LLM reasoning processes (Zhang et al., 2023b; Wu et al., 2022; Sel et al., 2023; Ning et al., 2023; Zhou et al., 2022; Zhang et al., 2023a). Tree of thoughts (ToTs) (Yao et al., 2023) models the reasoning process as a tree and employs DFS or BFS search to explore sequential thoughts. Tree-of-Thought by (Long, 2023) models the thoughts as trees but relies on a rule-based verifier to determine if a thought is valid and performs refining or backtracking based on a controller. Graph of Thoughts (Besta et al., 2023) models the process as a directed graph and defines 3 types of transformations at a node in the graph: aggregation, refining, and generation. These frameworks have better control over LLM’s intermediate steps, but they are not well-designed to consider LLM workflows with external tools and environments. StateFlow considers external feedback and also allows the design of complex patterns for more difficult tasks, where any state can be connected with the definition of state transitions. The step-wise search from ToTs (Yao et al., 2023) can easily be applied to StateFlow. When we call LLM in a state, we can generate several responses and employ an evaluator to select the best one.
LLMs have been used to interact with environments (Deng et al., 2023; Yao et al., 2022a; Shridhar et al., 2020) and tools (Paranjape et al., 2023; Gou et al., 2023; Schick et al., 2023; Gao et al., 2023; Yang et al., 2023b; Zhang et al., 2024). ReAct (Yao et al., 2022b) uses a few-shot prompting strategy that generates the next action based on past actions, and observations, which have been proven effective. Follow-up works employ iterative refining to improve from previous trials (Sun et al., 2024; Prasad et al., 2023). Reflexion (Shinn et al., 2023) and Self-Refine (Madaan et al., 2023) generate reflections from the past to improve future trials. In parallel to our work, Liu & Shuai (2023) adapts a state machine to record and learn from past trajectories. Extending from ToTs and RAP (Hao et al., 2023)(Zhou et al., 2023) proposes an LLM-based tree search incorporating reflection and feedback from the environment. These methods incur additional costs from the interactions or searching processes. StateFlow is orthogonal to these methods, and some of them can be used on top of StateFlow to further improve performance.
@@ A.1 Experiment Details
See Table 14 and 15 for the instructions used for SQL StateFlow model and Table 16 and 17 for instructions used for Bash. Each instruction is constructed with a uniform prompt that introduces the overall environment and a specific prompt that instructs the model to act. We follow the InterCode benchmark to use the chat-based models as text completions. All the instructions are put in the system message of the model. The task is entered as the user message, and new responses from the model and environment will be appended to that user message, following a completion manner.
Code is available at https://github.com/kaebvcidn/StateFlow.
@@ A.3 Additional Results
StateFlow 69.34 0.7223 6.11 1.89 36.0 2700 261
SF_Chat 70.41 0.7329 5.84 1.20 49.2 3907 283
Table 5: Results of the Intercode SQL with GPT-3.5 and GPT-4 with all metrics. We also include SF_Chat, an alternative of StateFlow, and another baseline Try Again with an oracle setting(). Best metrics of each model is in Bold. Second-best is Underlined.
@@
Table 6: Results of the Intercode Bash with GPT-3.5 and GPT-4 with all metrics. We also include SF_Chat, an alternative of StateFlow, and another baseline Try Again with an oracle setting().
@@ B.3 Additional Experiments
Head of system message
Interact with a self.setting system using self.language queries to answer a question.
## RESPONSE FORMAT
For action, put your SQL command in the execute[] block. You should only give one command to execute per turn. Reply with the following template (<<<>>> is the field description, replace it with your own response):
Thought: $<$your thought$>$
Action: execute[$<$your command$>$]
Observe ## Instructions
@@
Follow the RESPONSE FORMAT to give your thought and action.
@@
## Examples Thought: A ‘transcripts‘ table exists, but it doesn’t have the ‘release_date‘ column I came up with. I should find out what columns are available. Thought: The ‘friends‘ table has two ids. I should check if the ‘highschooler‘ table has a name associated with an ID. Thought: The ‘contestants‘ is a table, it is not a column in ‘people‘. I need to check the ‘contestants‘ table to see how to get the contestant names. Thought: I get a single number that is the number of likes that the high schooler Kyle has. This should be the answer.
Follow the RESPONSE FORMAT to give your thought and action.
@@
Table 14: Instructions of StateFlow for SQL task (part 1).
Verify
## Examples Thought: The output displays the contestant names and also contestant count. Although the count is used for sorting, it should not be displayed in output. I should modify my query to only select the contestant names.
@@
Reply with the following template (<<<>>> is the field description, replace it with your own response, ”—” is the ”or” operation):
Action: execute[$<$your new command$>$] — submit
@@
Head of system message Interact with a self.setting system using self.language queries to answer a question.
Use interleaving Thought, Action, Observation steps.
Each time, give a thought and an action, and then observe the result of the action.
For action, put your self.language command in the execute[] block. You should only give one command to execute per turn.
## OUTPUT DESCRIPTION
The user will provide the output of the action with this format:
Observation: $<$string$>$
## RESPONSE FORMAT
Reply with the following template (<<<>>> is the field description):
Thought: $<$your thought in one sentence$>$
Action: execute[$<$your command$>$]
@@
Please carefully review last command to determine why last observation didn’t return any output.
- If last action is a command that won’t return any output (e.g., ’touch’), please write an observation command to check last action.
- If it is possible for last action to return output (such as ’find’), please check if there is any mistake in your command, revise and modify your command as needed.
@@
Thought: Seems like this did not produce any results. I may have to use another command instead of cut to pick out the hashes from the output string
@@
Thought: I should check that the file was created successfully.
Action: execute[ls testbed/]
@@
For action, put your self.language command in the execute[] block. Only give one command per turn.
Thought: $<$your thought on whether the task is solved and last output answers the question.$>$
Action: execute[$<$your command$<$]

### v3

@@ 1 Introduction
LLMs have increasingly been employed to solve complex, multi-step tasks. Specifically, they have been applied to tasks that require interactions with environments (Yang et al., 2023a; Yao et al., 2022a; Shridhar et al., 2020; Zelikman et al., 2022) and those tasks that can benefit from utilizing tools such as web search and code execution (Mialon et al., 2023; Wu et al., 2023b; Wang et al., 2023c). In approaching these tasks, there is typically a desired workflow, or plan of actions based on heuristics that could improve the efficiency of task solving (Kim et al., 2023; Wu et al., 2023a). A common practice in the context of LLMs, such as ReAct (Yao et al., 2022b) and the vast customization of GPTs, is to write a single prompt that instructs the models to follow a desired procedure to solve the task (Dohan et al., 2022). The LLM is called iteratively with the same instruction, along with previous actions and feedback from tools/environments. This relies on LLMs’ innate capability to determine the current task-solving status and perform subsequent actions autonomously. Despite the impressive abilities of LLMs, it is still unrealistic to expect LLMs to always make the correct judgment of the status of current progress. It is also almost impossible to reliably track these judgments and their decisions of subsequent action trajectory. Given these considerations, we pose the research question: How can we exert more precise control and guidance over LLMs?
@@ 4.1 InterCode Benchmark
SRnormal-$\uparrow$ Turnsnormal-$\downarrow$ Errornormal-$\downarrow$ Costnormal-$\downarrow$ SRnormal-$\uparrow$ Turns $\downarrow$ Errornormal-$\downarrow$ Cost $\downarrow$
@@
SRnormal-$\uparrow$ Turnsnormal-$\downarrow$ Errornormal-$\downarrow$ Costnormal-$\downarrow$ SRnormal-$\uparrow$ Turns $\downarrow$ Errornormal-$\downarrow$ Cost $\downarrow$
@@
SR Turns Error Cost % $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$ StateFlow 63.73 5.67 6.82 3.82 No_Verify 62.28 5.18 5.96 3.68 No_Error 58.80 5.72 11.6 4.05 No_Obsrve 57.83 6.00 17.0 4.64
@@ 4.2 ALFWorld
Pick Clean Heat Cool Look Pick 2 Allnormal-$\uparrow$ Cost \$ $\downarrow$
@@ 5 Related work
Different prompting frameworks have been proposed to enhance LLM reasoning processes (Zhang et al., 2023b; Wu et al., 2022; Sel et al., 2023; Ning et al., 2023; Zhou et al., 2022; Zhang et al., 2023a; Zelikman et al., 2023). Tree of thoughts (ToTs) (Yao et al., 2023) models the reasoning process as a tree and employs DFS or BFS search to explore sequential thoughts. Tree-of-Thought by (Long, 2023) models the thoughts as trees but relies on a rule-based verifier to determine if a thought is valid and performs refining or backtracking based on a controller. Graph of Thoughts (Besta et al., 2023) models the process as a directed graph and defines 3 types of transformations at a node in the graph: aggregation, refining, and generation. These frameworks have better control over LLM’s intermediate steps, but they are not well-designed to consider LLM workflows with external tools and environments. StateFlow considers external feedback and also allows the design of complex patterns for more difficult tasks, where any state can be connected with the definition of state transitions. The step-wise search from ToTs (Yao et al., 2023) can easily be applied to StateFlow. When we call LLM in a state, we can generate several responses and employ an evaluator to select the best one.
LLMs have been used to interact with environments (Deng et al., 2023; Yao et al., 2022a; Shridhar et al., 2020) and tools (Paranjape et al., 2023; Gou et al., 2023; Schick et al., 2023; Gao et al., 2023; Yang et al., 2023b; Zhang et al., 2024; Zou et al., 2024). ReAct (Yao et al., 2022b) uses a few-shot prompting strategy that generates the next action based on past actions, and observations, which have been proven effective. Follow-up works employ iterative refining to improve from previous trials (Sun et al., 2024; Prasad et al., 2023). Reflexion (Shinn et al., 2023) and Self-Refine (Madaan et al., 2023) generate reflections from the past to improve future trials. In parallel to our work, Liu & Shuai (2023) adapts a state machine to record and learn from past trajectories. Extending from ToTs and RAP (Hao et al., 2023)(Zhou et al., 2023) proposes an LLM-based tree search incorporating reflection and feedback from the environment. These methods incur additional costs from the interactions or searching processes. StateFlow is orthogonal to these methods, and some of them can be used on top of StateFlow to further improve performance.
@@ A.1 Experiment Details
See Table 14 and 15 for the instructions used for SQL StateFlow model and Table 16 and 17 for instructions used for Bash. A uniform prompt that introduces the overall environment is put in the system message, and a specific prompt is put in the head of the user message.
Code is available at https://github.com/kevin666aa/StateFlow.
@@ A.3 Additional Results
StateFlow 69.34 0.7223 5.11 1.89 36.0 2700 261
SF_Chat 70.41 0.7329 4.84 1.20 49.2 3907 283
Table 5: Results of the Intercode SQL with GPT-3.5 and GPT-4 with all metrics. We also include SF_Chat, an alternative of StateFlow, and another baseline Try Again with an oracle $*{}^{*}$). Best metrics of each model is in Bold. Second-best is Underlined.
@@
Table 6: Results of the Intercode Bash with GPT-3.5 and GPT-4 with all metrics. We also include SF_Chat, an alternative of StateFlow, and another baseline Try Again with an oracle $*{}^{*}$).
@@ B.3 Additional Experiments
System message
Interact with a MySQL Database system using SQL queries to answer a question.
Observe
## Instructions
@@
## RESPONSE FORMAT
For action, put your SQL command in the execute[] block.
Thought: $<$your thought on which table(s) is/are relevant in one short sentence$>$
Action: execute[$<$your command$>$]
@@
## Examples
Thought: A ‘transcripts‘ table exists, but it doesn’t have the ‘release_date‘ column I came up with. I should find out what columns are available.
Thought: The ‘friends‘ table has two ids. I should check if the ‘highschooler‘ table has a name associated with an ID.
Thought: The ‘contestants‘ is a table, it is not a column in ‘people‘. I need to check the ‘contestants‘ table to see how to get the contestant names.
Thought: I get a single number that is the number of likes that the high schooler Kyle has. This should be the answer.
## RESPONSE FORMAT
For action, put your SQL command in the execute[] block. You should only give one command to execute per turn.
Thought: $<$your thought on why this query is error and whether you should gather more information or fix the error in one sentence$>$
Action: execute[$<$your command$>$]
Verify
## Examples
Thought: The output displays the contestant names and also contestant count. Although the count is used for sorting, it should not be displayed in output. I should modify my query to only select the contestant names.
## RESPONSE FORMAT
Reply with the following template (<<<>>> is the field description, replace it with your own response, ”||||” is the ”or” operation):
Action: execute[$<$your new command$>$] |||| submit
Table 14: Instructions of StateFlow for SQL task (part 1).
@@
For action, put your SQL command in the execute[] block.
Thought: $<$your thought on constructing command to answer the query exactly$>$
Action: execute[$<$your command$>$]
@@
System message
Interact with a Bourne Shell system using BASH queries to answer a question. Follow the user’s instructions to solve the problem.
@@
## RESPONSE FORMAT
For action, put your BASH command in the execute[] block. Only give one command per turn.
Thought: $<$your thought in one sentence$>$
Action: execute[$<$your command$>$]
@@
## RESPONSE FORMAT
Reply with the following template (<<<>>> is the field description):
Thought: $<$your thought in one sentence$>$
Action: execute[$<$your command$>$]
@@
Carefully check if the question is answered.
- Please check if the desired tasks have been performed.
- If the question also asks for output, please check your last command and output, and make sure the output is in the desired format, and doesn’t contain any extra fields.
- If the desired tasks have been performed, please submit the query with this ”Action: submit” command.
@@
Thought: This gives me storage information for every folder under the workspace directory, but I only need the storage for just the ‘workspace/‘ directory. The ‘-s‘ option should help with this.
Action: execute[du -sh /workspace]
Thought: This shows the output hashes and they have the same values, indicating that these files are duplicates. However, the file names are also shown, which are not needed.
@@
Thought: This shows me too much information, I only need the total number of lines. I should add up the lines together and output a single number.
Thought: The hello.txt file has been created successfully in the testbed/ directory, and it contains the Hello World text. I can submit.
Action: submit
Please follow this RESPONSE FORMAT to give your thought and action.
@@
For action, put your {self.language} command in the execute[] block. Only give one command per turn. If the question is solved, your action should be ”Action: submit”.
Reply with the following template (<<<>>> is the field description, replace it with your own response, ”||||” is the ”or” operation):
Thought: $<$your thought on whether the question is answered in one sentence$>$
Action: execute[$<$your new command$>$] — submit